there are three prizes. the value of the second prize is five-sixths the value of the first prize. the third prize is four-fifths that of the second prize.
I have to express the total value of the three prizes as an algebraic expression.

Answer by stanbon(48568)

there are three prizes. the value of the second prize is five-sixths the value of the first prize. the third prize is four-fifths that of the second prize.
I have to express the total value of the three prizes as an algebraic expression.
-----------------
Let 1st be "x".
2nd is (5/6)x
3rd is (4/5)(5/6)x = (2/3)x
----------------
Total value = x + (5/6)x + (4/6)x
Total value = (15/6)x or 2 1/2 x
=================================
Cheers,
Stan H.